The ionization constant of HA is 1.0 × 10−4 . What must be the ratio of [A−] to [HA] for the preparation of a buffer solution with a pH of 4.87?


1
Expert's answer
2021-03-10T06:57:53-0500

In a buffer solution where HA is the weak acid/base.

HA(aq)⇌H+(aq)+A−(aq)

Given data:-

Ionisation constant (Ka)=1.0*10^-4

We know PKa= -logKa

So, PKa= -log(1.0*10^-4)=4

PH=4.87

According to Henderson equation we get-

PH=PKa +log(A-/HA)

4.87=4+log(A-/HA)

log(A-/HA)=4.87-4

log(A-/HA)=.87

A-/HA=10^(. 87)

A-/HA=7.41

Hence the ratio of [A−] to [HA] for the preparation of a buffer solution is 7.41